OpenStudy (dbzfan836):

What is the relationship between mass and momentum? Be specific as you describe the relationship between mass and linear momentum as well as how mass affects the moment of inertia in angular momentum.

OpenStudy (dbzfan836):

do u have an answer?

OpenStudy (abhisar):

Hello @Dbzfan836 !

OpenStudy (abhisar):

Momentum (P) = Mass \(\times\) Velocity

OpenStudy (dbzfan836):

I havent seen you for awhile abishar

OpenStudy (dbzfan836):

*@Abhisar

OpenStudy (abhisar):

\(\Huge{\overset{\frown}{\normalsize \left( \begin{matrix} \Large\cdot \quad \cdot\\ \cdot\\ \huge \smile \end{matrix} \right)}}\normalsize \\ \;/\quad \;\;\quad \backslash\)

OpenStudy (dbzfan836):

well, how does it affect interia?

OpenStudy (abhisar):

Inertia or moment of inertia ?

OpenStudy (dbzfan836):

both

OpenStudy (abhisar):

For any object Inertia \(\propto\) Mass

OpenStudy (abhisar):

Moment of Inertia (I) = M\(\sf R^2\), where M= mass and R=Distance between the axis and rotation. So again I \(\propto\)Mass

OpenStudy (dbzfan836):

oh, thank you!

OpenStudy (abhisar):

\(\rlap{\color{blue}{\huge\bigstar}\huge \color{blue}{ \text{You're Most Welcome! }}\color{blue}\bigstar}{\; \color{aqua}{\huge\bigstar}\huge \color{aqua}{\text{You're Most Welcome! }}\color{aqua}\bigstar}\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{green}{\huge\ddot\smile}\color{blue}{\huge\ddot\smile}\color{pink}{\huge\ddot\smile}\color{red}{\huge\ddot\smile}\color{yellow}{\huge\ddot\smile}\)